# Enhancing Pre-Trained Decision Transformers with Prompt-Tuning Bandits

### v2

@@ 4.2 Scalable and sample-efficient bandit architecture
5:        Make PDT input (Eq. 2): $\mathbf{x}_{k}=\rho_{k}\odot\omega_{K:t}$
@@
Algorithm 2 CMAB Prompt Selection
@@
2:     Predict reward $\mathbf{\hat{y}}=[\phi_{1}(\tilde{\tau}),\dots,\phi_{J}(\tilde{\tau})]$
@@
Algorithm 3 CMAB Update
@@
4.3 Arm features
Learning reward models $\phi_{j}:\tilde{\tau}\to\mathbb{R}$ for “raw” trajectory segments becomes impractical for MDPs with large state spaces, such as those involving pixel-space observations and high-dimensional actions. The key issue lies in the input size of these reward models, which scales as $H\times(|\mathcal{S}|+|\mathcal{A}|+1)$, growing linearly with $\mathcal{S},\mathcal{A}$, and $H$. This limitation can be addressed by integrating a feature extractor $\Psi:{\tilde{\tau}}\to\mathbb{R}^{d}$ that embeds raw trajectory segments from the input space into a latent feature space. Such a feature extractor can be pre-trained using various methods on the demonstration dataset, enabling the reward models to operate more efficiently in the resulting embeddings’ space.
Alternatively, the pre-trained PDT model itself can be leveraged as a feature extractor by utilizing the hidden representation of prompt tokens as the embedding for the prompt. This approach not only mitigates the scaling issue but also aligns with the inductive biases of the pre-trained PDT, which is expected to encode meaningful task-relevant information. The reward models then operate on these fixed-size embeddings rather than raw, flattened segments, significantly improving scalability.
5 Experiments
In this section, we evaluate the proposed prompt-tuning method, focusing on whether inference-time performance gains can be attributed to prompt tuning. Additionally, we analyze its impact on prompt selection, assess improvements in OOD generalization, examine robustness to prompt dataset quality, and investigate performance in high-dimensional state and action spaces.
5.1 Environments, baselines, and implementation
(a) Multiple tasks.
(b) Exemplary trajectories.
Figure 2: Our Sparse 2D point environment. Left: Multiple tasks can be induced through the task parameters $r$ and $\alpha$, corresponding to radius and angle. Right: Expert and novice trajectories for a particular task.
Environments. While most offline RL and DT-based studies have predominantly used MuJoCo (Todorov et al., 2012) benchmarks as test beds, the tasks in this benchmark often lack the features required for a robust multi-task generalization and prompt-tuning study. To address this limitation, experimental results are reported not only on MuJoCo but also on a newly designed “Sparse 2D point” environment specifically tailored to better evaluate multi-task generalization.
@@ 5.1 Environments, baselines, and implementation
Sparse 2D point: This task involves mixed control of a planar point agent, where the state representation consists of the agent’s current 2D coordinates. The agent always begins at $(0,0)$ with the objective of reaching a specified goal coordinate. It has two continuous actions for movement across the plane and a binary stop action. The task requires the agent to navigate to the goal coordinate and appropriately select the stop action. A sparse reward is provided upon selecting the stop action, proportional to the agent’s distance from the goal. When stop is selected within close proximity of the goal, the environment provides a reward bonus of 10, which is discounted based on the number of wasted steps. Tasks are distinguished by varying goal coordinates, distributed on circles with different radii and angles, as depicted in Figure 2(a). While the goal is not explicitly part of the state, it is implicitly encoded through the reward function, ensuring that each task remains a fully observable MDP.
@@
MuJoCo Half Cheetah: As introduced by Xu et al. (2022), this task involves continuous joint control of a planar “half-cheetah” agent. The action space and state space have dimensionalities of 7 and 20, respectively. Different tasks are defined by varying target velocities for the agent, with the reward being proportional to the discrepancy between the agent’s actual and target velocities. Although the target velocity is not explicitly included in the state, it is implicitly encoded through the reward function, ensuring that each task corresponds to a fully observable MDP.
Baselines. Our proposed method is evaluated against the following RL baselines: (i) an optimal policy oracle, (ii) a DT trained on multiple tasks without additional enhancements, (iii) and a vanilla PDT, which samples prompts uniformly at random, (iv) and a Gaussian perturbation-based algorithm, where we effectively perform hill climbing in the prompt space, similarly to Hu et al. (2023).
Implementation details. All experiments were conducted on an instance equipped with NVIDIA T4 GPU with 8GB of memory, utilizing the PyTorch library (Paszke et al., 2019). Details of all hyperparameter configurations are provided in Appendix B.
Offline reinforcement learning dataset For each training task $i$, PDT requires a dataset of trajectories $\mathcal{T}_{i}$ and a dataset of expert demonstrations $\mathcal{P}_{i}$. For the Mujoco Half Cheetah tasks, we use datasets from Xu et al. (2022) comprising 40 tasks, each corresponding to a distinct velocity. Of these, five tasks $\{2,7,15,23,26\}$ are reserved for evaluation, and the other 35 tasks are used for training.
To generate datasets for the Sparse 2D point environment, we employ PPO (Schulman et al., 2017) on 60 tasks with three discrete radii and 20 discrete angles. The first 48 tasks are used for training, and the remaining 12 are reserved for testing generalization. For more task details see Sec. A in the Appendix. PPO is run for 1M steps on each training task, and the resulting trajectories are stored as $\mathcal{D}_{i}$. The highest-return trajectories are selected to form $\mathcal{P}_{i}$. To reduce the computational cost of computing $\arg\max$ over all possible prompt segments, $\mathcal{P}_{i}$ is limited to 10 randomly sampled high-return trajectories.
5.2 Results and Analysis
Figure 3: Spatio-temporal visualization of prompt selection across tasks. Each dot represents the mean segment coordinates, colored by performance. The red diamond marks the starting state, while the red stars indicate the goal coordinates of tasks used in the experiment. The shaded region indicates the holdout test tasks.
Method $J=1$ $J=2$ $J=4$
PDT, no tuning 0.0 $\pm$ 2.1 6.3 $\pm$ 0.8 8.3 $\pm$ 0.6
$\epsilon$-GreedyΨ 9.0 $\pm$ 0.6 9.4 $\pm$ 0.3 9.6 $\pm$ 0.2
UCBΨ 9.2 $\pm$ 0.5 8.8 $\pm$ 0.9 8.8 $\pm$ 0.9
$\epsilon$-Greedy, raw 8.9 $\pm$ 0.5 9.5 $\pm$ 0.3 8.9 $\pm$ 0.7
UCB, raw 9.4 $\pm$ 0.5 9.5 $\pm$ 0.3 9.3 $\pm$ 0.4
Gaussian perturbation 5.8 $\pm$ 3.8 7.9 $\pm$ 1.6 6.2 $\pm$ 4.0
Standard DT -64.3 $\pm$ 24.1
Table 1: Inference time performance in the Sparse 2D point environment. $\Psi$ denotes that the bandit’s reward model operates on encoded trajectory segments; otherwise, it processes non-encoded trajectory segments. Results are averaged across training tasks, three random seeds, and the last 50 rollouts.
Method $J=1,H=5$ $J=2,H=20$
PDT, no tuning -44.75 $\pm$ 0.91 -42.68 $\pm$ 1.3
[…]
$\epsilon$-greedy CMAB -76.50 $\pm$ 6.68 -58.76 $\pm$ 8.21
Table 2: Inference time performance in the Mujoco Half Cheetah environment. Columns represent models with increasing numbers of trajectory segments ($J$) and segment lengths ($H$) in the prompt. Results are averaged over multiple tasks, seeds, and rollouts.
Does bandit-based prompt-tuning improve the inference-time performance of a frozen PDT backbone?
To investigate this question empirically, we compare the inference-time performance of a standard PDT model trained to convergence, sampling demonstrations randomly from $\mathcal{P}_{i}$, with the same model enhanced by our bandit-based prompt selection method. We conduct 250 online rollouts on the training tasks with the largest radius, both with and without prompt tuning and report the mean performance from the final 50 rollouts, averaged over three seeds. Results for the Sparse 2D point environment are presented in Table 1.
The results show that the proposed method consistently enhances the PDT backbone’s performance, achieving optimal returns. This effect is most prominent in the $J=1$ setting, as larger $J$ (i.e., more segments and tokens in the prompt) naturally improves PDT’s performance, even without prompt-tuning, reducing the performance margin for tuning. This implies that, for the Sparse 2D Point environment, randomly sampling large prompts suffices for finding tokens that enable the PDT to identify the target task. We observe no significant differences between $\epsilon$-greedy and UCB-based exploration strategies or between training the bandit’s reward models using “raw” trajectory segments compared to Transformer-based representations, as described in Section 4.3.
Compared to the Gaussian perturbation baseline, bandit-based tuning achieves consistently higher returns with substantially lower variance. Notably, Gaussian perturbation reduces performance when $J>1$, likely because it introduces noise across the entire prompt, including segments that are already effective. In contrast, our bandit approach is more efficient and robust, since it models each segment as an independent arm. This allows it to retain optimal segments while continuing to explore others, reliably boosting PDT to optimal performance even for $J>1$. The standard DT baseline further highlights the importance of prompts in distinguishing tasks. Additional plots are provided in Appendix C.1.
Results for the higher-dimensional MuJoCo Half Cheetah task are presented in Table 2. In this more complex environment, the proposed method provides substantial performance improvements over the baseline pre-trained model. As $J$ and $H$ increase, the gains from the bandit approach become even more pronounced.
[…]
Which prompts does bandit-tuning select?
We qualitatively analyze how the bandit’s prompt selection strategy evolves over time and how it learns to identify high-performance segments. For this analysis, we use PDT with one segment per prompt ($J=1$) and $\epsilon$-greedy exploration, where $\epsilon$ is annealed from 1 to 0 over the first 30 rounds.
Figure 3 illustrates the prompt segments selected by the bandit over an increasing number of rounds for 20 tasks corresponding to different angles. The segments are represented by their mean spatial coordinates. During the initial rounds, the bandit explores by uniformly sampling segments from the prompt datasets $\mathcal{P}_{i}$, encountering both low- and high-performing prompts. In this Sparse 2D point environment, low-performance prompts are associated with segments closer to the starting state at the circle’s center, as these fail to clearly indicate where the agent should stop. Over time, the bandit exploits its accumulated knowledge to disregard these low-performance prompts and instead selects segments nearer to the goal coordinates, which are more informative for task identification.
Can prompt-tuning exploit non-expert datasets?
@@ 5.2 Results and Analysis
While the PDT model with a random prompt selection strategy relies on access to expert demonstrations, our method demonstrates greater robustness to the quality of demonstration data. To examine this, we generate additional prompt datasets $\{\mathcal{P}_{i}^{\%j}\},j\in\{0,10,\dots,100\}$, which combine $j\%$ of expert data along with novice demonstrations, selected from trajectories in the bottom 5th percentile of task returns (see Fig. 2(b)).
@@
Expert percentage $j\%$ No tuning $\epsilon$-greedy UCB
@@
Does prompt-tuning improve OOD task generalization?
To assess the benefit of prompt-tuning on out-of-distribution (OOD) tasks, we evaluate the pre-trained PDT model on OOD tasks within the Sparse 2D point environment, where goal locations lie on the largest circle with angles exceeding those in training, as depicted in Figure 3. We perform 250 rollouts, calculate the mean over the final 50, and average the results across three seeds. Results in Table 4 show that prompt-tuning consistently improves performance, achieving gains comparable to the in-distribution setting, though not always reaching optimal levels.
Further OOD generalization results for the Mujoco Half Cheetah environment are presented in Table 5. Using the same evaluation protocol (multiple tasks, 250 rollouts with the mean of the final 10 averaged across three seeds), we observe that the bandit-based prompt-tuning method significantly enhances PDT’s generalization to OOD tasks. Notably, this improvement becomes increasingly evident with larger prompt sequences (i.e., increasing $J$ and $H$), further underscoring the efficacy of this approach.
Notably, fine-tuning the PDT backbone exclusively on target task data for 250 epochs results in a performance decline across both tested environments. This degradation is hypothesized to stem from the large size of the Transformer backbone, the absence of regularization provided by diverse multi-task training data, and a misalignment between offline action loss and online task performance. These factors likely lead to overfitting on the fine-tuning dataset without yielding improved online returns. Supporting plots are provided in Appendix C.3.
Method Pre-trained $\theta$ Fine-tuned $\theta$
PDT, no tuning -6.7 $\pm$ 7.0 -28.3 $\pm$ 22.7
With $\epsilon$-greedy bandit -2.4 $\pm$ 8.7 -21.5 $\pm$ 22.3
Table 4: Performance on OOD tasks in Sparse 2D point environment. Columns differentiate between the pre-trained and fine-tuned PDT parameter.
Method $J=1,H=5$ $J=2,H=20$
PDT, no tuning -36.83 $\pm$ 2.58 -35.04 $\pm$ 3.60
PDT, fine-tuning -71.51 $\pm$ 15.85 -61.57 $\pm$ 8.09
[…]
Table 5: Performance on OOD tasks in the Mujoco Half Cheetah environment. Columns represent models with increasing numbers of trajectory segments ($J$) and segment lengths ($H$) in the prompt.
6 Conclusion
Limitation and broader impact
A key limitation of the proposed method is its high variance, stemming from the assumption that each reward model treats its corresponding segment as the sole factor influencing performance, and ignoring correlations between segments. This simplification opens avenues for future work aimed at disentangling these dependencies to improve model robustness and accuracy. Another challenge arises when the size of the expert demonstration dataset becomes prohibitively large. Iterating over the entire dataset becomes computationally expensive, especially given the combinatorial growth factor associated with the number of prompt segments, $J$. A potential solution could involve learning a sampler to identify promising segments, which could then be paired with reward models for further refinement. These areas highlight opportunities for advancing the methodology to address scalability and efficiency.
Appendix
Appendix A Sparse 2D point environment details
Our Sparse 2D point environment offers a continuous task space, parameterized by the angle and radius, for arbitrary goal locations on the 2D plane. We discretize the task space by using three discrete radii, $(0.9,1.9,2.9)$, and 20 discrete angles $(0.1\cdot\pi,0.2\cdot\pi,\dots,1.9\cdot\pi)$, instead of sampling continuous task parameters. This is primarily done to separate datasets for different tasks, since PDT requires expert demonstration for each training task. To separate these $3\cdot 20=60$ tasks into the training set $\mathcal{T}^{\text{train}}$ and testing set $\mathcal{T}^{\text{test}}$, all tasks with an angles greater than $1.5\cdot\pi$ (independently of the radius) are treated as testing task and are not part of the training set. This split yields 48 training tasks, and 12 testing tasks. Spatially, the test set is indicated by the shaded are in Figure 3.
Appendix B Training details
Details of the hyperparameters for the DT and PDT models are provided in Table 7, while those for the bandit model are listed in Table 7. Game-specific hyperparameters are reported separately in Table 8. The implementations of DT and PDT were adopted from Barhate (2022) with minimal modifications to integrate seamlessly with the CMAB prompt-tuning framework. For the Gaussian perturbation baseline, we sample an initial prompt randomly from the expert demonstration dataset for the target task, then apply Gaussian noise and perform hill climbing. We linearly anneal the scale of the exploration noise from 1.1 to 0.1 over the course of the 250 online rollouts.
Our source code is publicly available at https://retracted-during-review.
Hyperparameter Value
Number of transformer blocks 3
Number of attention heads 1
Embedding dimension 128
Transformer activation function GELU
MLP activation function ReLU
MLP hidden layers 2
MLP width 128
Batch size (per task) 8
Learning rate 1e-4
Learning rate decay weight 1e-4
Table 6: Hyperparameter values for PDT and DT models.
Hyperparameter Value
Batch size All data
Learning rate 1e-3
Evaluation trials 250
$\epsilon$ in $\epsilon$-greedy 0.1
$c$ (exploration parameter) in UCB 3
Reward model type MLP
MLP hidden layers 2
MLP width 16
MLP activation function ReLU
Table 7: Hyperparameters of Contextual Multi-armed Bandit
Environment Target Return Prompt Length $(J\times H)$ Context Length
2D-Point-Sparse 10 $(1\times 3),(2\times 3),(4\times 3)$ 5
Cheetah-vel 0 $(1\times 5),(2\times 20)$ 20
Table 8: Environment-specific hyperparameters of DT and PDT.
Appendix C Appendix: Additional plots
C.1 Online rollouts in Sparse 2D point
(a) Online performance of standard PDT, without prompt-tuning.
(d) Online performance with prompt-tuning, using $\epsilon$-greedy exploration and raw trajectory segments instead of transformer features for the bandit’s reward models.
(e) Online performance with prompt-tuning, using UCB exploration and raw trajectory segments instead of transformer features for the bandit’s reward models.
Figure 4: Plots of data used in Table 1, averaged over three seeds and all training tasks. The dashed line marks the optimal return at +10.
C.2 Prompt quality experiment for the Sparse 2D point environment
@@
C.3 Out-of-distribution finetuning
(a) Offline action loss
(b) Online task return
Figure 6: Fine-tuning of the PDT backbone on individual OOD tasks from our Sparse 2D point environment indicating a disconnect between the offline action loss and online task performance. Results are averaged for all fine-tuning tasks and PDT backbones trained with three different random seeds. The shaded area indicates one standard deviation around the mean.

### v3

@@ Appendix A Prompt-Tuning Algorithm
5:   Make PDT input (Equation 3): $\mathbf{x}_{k}=\rho_{k}\odot\omega_{L:t}$
@@
Algorithm 2 Prompt Selection
@@
[…]
@@
Algorithm 3 Bandit Update
@@
Appendix B Environment Details
@@
Sparse 2D point: This environment involves mixed control of a planar point agent, where the state representation consists of the agent’s current 2D coordinates. The agent always begins at $(0,0)$ with the objective of reaching a (hidden) goal coordinate. It has two continuous actions for movement across the plane and a binary stop action. The task requires the agent to navigate to the goal coordinate and appropriately select the stop action. A sparse reward is provided upon selecting the stop action, proportional to the agent’s distance from the goal. When stop is selected within close proximity of the goal, the environment provides a reward bonus of 10, which is discounted based on the number of wasted steps. While the goal is not explicitly part of the state, it is implicitly encoded through the reward function, ensuring that each individual task remains a fully observable MDP.
The environment offers a continuous task space, parameterized by the angle and radius, for arbitrary goal locations on the 2D plane. We discretize the task space by using three discrete radii, $(0.9,1.9,2.9)$, and 20 discrete angles $(0.0\cdot\pi,0.1\cdot\pi,\dots,1.9\cdot\pi)$, instead of sampling continuous task parameters. This is primarily done to separate datasets for different tasks, since PDT requires expert demonstration for each training task. To separate these $3\cdot 20=60$ tasks into the training set $\mathcal{T}^{\text{train}}$ and testing set $\mathcal{T}^{\text{test}}$, all tasks with an angles greater than $1.5\cdot\pi$ (independently of the radius) are treated as testing task and are not part of the training set. This split yields 48 training tasks, and 12 testing tasks. Spatially, the test set is indicated by the shaded are in Figure 3.
@@
MuJoCo Half Cheetah: This task involves continuous joint control of a planar half-cheetah agent. The state and action spaces have dimensions 20 and 7, respectively. Tasks vary by target velocity, and the reward is proportional to the deviation from this target. Although the target velocity is not part of the observable state, it is implicitly encoded via the reward, ensuring that each task defines a fully observable MDP. There are 40 tasks in total, 35 are used for training and 5 are used for testing. For more environment details see (Duan et al., 2016; Xu et al., 2022).
MuJoCo Ant: As in Xu et al. (2022), this task requires continuous control of an ant agent via eight joint actuators. The 27-dimensional state space includes positions and velocities of the agent’s body. Tasks involve moving in different directions, with reward based on velocity along the target heading. There are 50 tasks in total, 45 are used for training and 5 are used for testing. For more environment details see Duan et al. (2016); Xu et al. (2022).
Meta-World Pick-Place: Following Yu et al. (2020), this task involves Cartesian control of a simulated Sawyer robot’s end-effector. The 4-dimensional action space includes 3D position deltas and gripper torque. The 39-dimensional state space encodes gripper and object positions, and quaternions, but excludes the goal location. The benchmark consists of 50 task variations with different placement goals; 5 are held out during training for evaluation.
Appendix C Training Details
All experiments were conducted on an instance equipped with NVIDIA T4 GPU with 8GB of memory, utilizing the PyTorch library (Paszke et al., 2019). Pre-training the multi-task PDT backbone took approximately 5 hours. Performing 250 online rollouts and prompt tuning took approximately 30 to 120 minutes per task, depending on environment and algorithm.
Details of the hyperparameters for the DT and PDT models are provided in Table 4, while those for the bandit model are listed in Table 4. We used standard values for each hyperparameter instead of performing an expensive hyperparameter optimization. Environment-specific hyperparameters are reported separately in Table 5. The implementations of DT and PDT were adopted from Barhate (2022) with minimal modifications to integrate seamlessly with the CMAB prompt-tuning framework. The MuJoCo experiments were performed by integrating our prompt-tuning bandits into the official PDT repository Xu et al. (2022). For the hill-climbing and ZORankSGD (Hu et al., 2023) baselines, we sample an initial prompt randomly from the expert demonstration dataset for the target task, then apply Gaussian noise and perform hill climbing. To encourage convergence to a final prompt, we linearly anneal the scale of the exploration noise from 1.1 to 0.1 over the course of the 250 online rollouts.
Hyperparameter Value
Number of transformer blocks 3
Number of attention heads 1
Embedding dimension 128
Transformer activation function GELU
MLP activation function ReLU
MLP hidden layers 2
MLP width 128
Batch size (per task) 8
Learning rate 1e-4
Learning rate decay weight 1e-4
Optimizer AdamW
Table 3: Hyperparameter values for PDT and DT models.
Hyperparameter Value
Batch size All data
Learning rate 1e-3
Evaluation trials 250
$\epsilon$ in $\epsilon$-greedy 0.1
$c$ (exploration parameter) in UCB 3
Reward model type MLP
MLP hidden layers 2
MLP width 16
MLP activation function ReLU
Optimizer Adam
Table 4: Hyperparameters of Contextual Multi-armed Bandit
Environment Target Return Prompt Length $(J\times H)$ Context Length
2D-Point-Sparse 10 $(1\times 3),(2\times 3),(4\times 3)$ 5
Cheetah-vel 0 $(1\times 5),(2\times 20)$ 20
Table 5: Environment-specific hyperparameters of DT and PDT.
Appendix D Additional Results
D.1 Online Rollouts in Sparse 2D point
(a) Online performance of standard PDT, without prompt-tuning.
(d) Online performance with prompt-tuning, using $\epsilon$-greedy exploration and unencoded trajectory segments instead of transformer features for the bandit’s reward models.
(e) Online performance with prompt-tuning, using UCB exploration and unencoded trajectory segments instead of transformer features for the bandit’s reward models.
Figure 5: Plots of data used in Table 3(b). The dashed line marks the optimal return at +10, the shaded area corresponds to 1 standard deviation around the mean.
(a) Online performance after 250 episode with prompt-tuning, using ZORankSGD for prompt tuning.
(b) Online performance after 1500 episodes with prompt-tuning, using ZORankSGD for prompt tuning.
Figure 6: Plots of ZORankSGD prompt-tuning as reported in Table 3(b). The dashed line marks the optimal return at +10. The shaded area corresponds to one standard deviation around the mean. ZORankSGD requires the online evaluation of $m=5$ prompt perturbations to compute one prompt update. Results using a budget of 250 online rollouts (and 50 prompt improvement steps) are shown in subfig (a). Results using 1250 online rollouts (and 250 prompt improvements) are shown in subfig (b).
D.2 Prompt Quality Experiment for the Sparse 2D point Environment
@@
Can prompt-tuning exploit non-expert datasets? While the PDT model with a random prompt selection strategy relies on access to expert demonstrations, our method demonstrates greater robustness to the quality of demonstration data. To examine this, we generate additional prompt datasets $\{\mathcal{P}_{i}^{\%j}\},j\in\{0,10,\dots,100\}$, which combine $j\%$ of expert data along with novice demonstrations, selected from trajectories in the bottom 5th percentile of task returns (see Figure 2).
@@
Expert percentage $j\%$
No tuning $\epsilon$-greedy UCB
@@
Appendix E Segment Independence Attention Analysis
The key assumption based on which we design our structured bandit architecture is that prompt segments contribute independently and additively to prompt informativeness. This assumption implies that PDT can identify the target tasks by attending to key $(\hat{r},\mathbf{s},\mathbf{a})$-pairs in the prompt, as opposed to deriving task identity from the interaction effects between prompt segments, or by attending to the prompt globally. This assumption is crucial, because it allows us to decompose the prompt tuning problem into a bandit that maintains $J$ reward models, one for each prompt segment, which results in a reduction of the size of the exploration problem from combinatorial in $J$ to linear in $J$. The basis for the assumption is that a) many MDPs can be characterized by their optimal state-action marginal (i.e. by key $(\mathbf{s},\mathbf{a})$-pairs) and b) that PDT’s pre-training paradigm discourages reliance on segment interaction or arrangement, since prompt segments are sampled uniformly at random during pre-training, not by enforcing order or other constraints between segments. This training setup implicitly encourages invariance to permutation and inter-segment dependencies, and thus we believe that this limits the influence of global prompt structure on the model’s behavior.
To provide additional support for this interpretation, we conduct an analysis of PDT’s attention weight on our 2D environment. Here, for each task, we construct prompts that contain exactly one segment that is highly informative for task identification, based on our domain knowledge, while the remaining J1J-1 segments carry little information for discriminating between tasks. We then performed rollouts using different permutations of these segments as prompts. The performance of PDT remained robust under those permutations with $6.27\pm 0.34$ return on average over tasks, prompts, and model configurations ($J=2,H=3$ and $J=4,H=3$), which implies prompt segment permutation invariance in PDT. We visualize representative PDT attention mask for different prompt permutations in Figure 8 to confirm qualitative that PDT indeed identifies tasks by attending mostly to the individual, informative prompt segment.
Furthermore, we quantify attention to each prompt segment by computing the mean of the token attention score of tokens in each prompt segment. We then sum the per-timestep mean segment attention scores, and find that in 99% of rollouts conducted during this analysis, PDT assigned highest attention to the informative segment. Additionally, we manually set attention weights for specific prompt segments to zero, and find that PDT’s performance remains stable when uninformative segments are masked out, while it degrades drastically when the informative segment is removed, as per Table 7.
Full prompt Only informative segment Only uninformative segments
$6.27\pm 0.34$ $5.88\pm 0.23$ $34.06\pm 4.49$
Table 7: PDT performance with different attention scopes.
These results suggest that PDT behavior is largely driven by the most informative segment, and that inter-segment interactions are negligible in practice, which supports and justifies our assumption and bandit architecture. However, we also observed instances where PDT attended to all prompt segments rather uniformly, despite J1J-1 segments not being useful for task discrimination as per our domain knowledge. We hypothesize that these instances were due to slight overfitting to the prompt dataset and see the design of regularization that prevents PDT from exploit such nuances in the prompt dataset as important future work.
(a) Prompt segment 1 is informative.
(b) Prompt segment 2 is informative.
(c) Prompt segment 3 is informative.
(d) Prompt segment 4 is informative.
Figure 8: Attention weights from the first transformer block (after softmax), darker values indicate higher attention scores. The first 20 rows have been cropped to improve readability. The blue horizontal lines indicate the indices of state tokens at times $t-H,t-H+1,\dots,t$ from which the action at the corresponds $t$ is predicted. It can be seen that PDT attends mostly to the informative prompt token, independently of it’s position or neighboring segments.
Appendix F Wall-clock inference time
Table 8 reports representative wall-clock inference times for different methods on the MuJoCo Ant environment. Comparing our bandit architecture with PDT-encoded prompt representations (denoted by Ψ) to its unencoded counterpart highlights the efficiency benefits of using the PDT as a feature extractor. When operating on PDT-encoded prompts, the Thompson Sampling bandit maintains low inference times even for large prompts $(J=2,H=20)$, since the PDT-encoded representation has a fixed dimension $\mathbb{R}^{d}$ determined by the PDT’s token embedding space, independent of the prompt length. In contrast, unencoded prompts lead to input sizes that scale as $H\times(|\mathcal{S}|+|\mathcal{A}|+1)$ for each segment, substantially increasing inference time due to the higher cost of updating the reward models.
Method MuJoCo Ant ($J=1,H=5$) MuJoCo Ant ($J=2$, H=20)
PDT (full-model fine-tuning) $\sim 28$m $\sim 38$m
Hill climbing $\sim 16$m $\sim 24$m
ZORankSGD $\sim 52$m 120\ \ $\sim$ 120m
TSΨ $\sim 28$m $\sim 34$m
TS $\sim 33$m 172\ \ $\sim$ 172m
Table 8: Wall-clock inference times with different methods.
Appendix G Proof of Theorem 4.1
Theorem.
Assume that the reward function $G\colon P^{J}\to\mathbb{R}$ for a prompt $\rho=(\tilde{\tau}_{1},\dots,\tilde{\tau}_{J})$ decomposes as the mean of $J$ independent reward models $\phi_{j}(\tilde{\tau}_{j})$:
$G(\rho)=\frac{1}{J}\sum_{j=1}^{J}\phi_{j}(\tilde{\tau}_{j})+h(\tilde{\tau}_{1},\dots,\tilde{\tau}_{J}),$
and that the interaction term is uniformly bounded by $|h(\tilde{\tau}_{1},\dots,\tilde{\tau}_{J})|\leq\varepsilon,\quad\forall\,\tilde{\tau}_{j}\in P.$ Let $\rho^{*}=(\tilde{\tau}^{*}_{1},\dots,\tilde{\tau}^{*}_{J})$ denote the optimal prompt, and suppose that for each slot $j$, a bandit algorithm guarantees a slot-specific regret
$\mathrm{Regret}_{j}(K)=\sum_{t=1}^{K}\mathbb{E}\left[\phi_{j}(\tilde{\tau}^{*}_{j})-\phi_{j}(\tilde{\tau}_{t,j})\right]$
over $K$ rounds. Then the cumulative regret after $K$ rounds is bounded as:
$\mathrm{Regret}(K)$ $\triangleq\sum_{t=1}^{K}\mathbb{E}\left[G(\rho^{*})-G(\rho_{t})\right]\leq\frac{1}{J}\sum_{j=1}^{J}\mathrm{Regret}_{j}(K)+2K\varepsilon.$
Proof.
We begin with the decomposition of the reward function. The instantaneous regret at round $t$ is
$G(\rho^{*})-G(\rho_{t})=\left(\frac{1}{J}\sum_{j=1}^{J}\phi_{j}(\tau_{j}^{*})+h(\tau_{1}^{*},\dots,\tau_{J}^{*})\right)-\left(\frac{1}{J}\sum_{j=1}^{J}\phi_{j}(\tau_{t,j})+h(\tau_{t,1},\dots,\tau_{t,J})\right).$
Rearranging terms yields:
$G(\rho^{*})-G(\rho_{t})=\frac{1}{J}\sum_{j=1}^{J}\Bigl(\phi_{j}(\tau_{j}^{*})-\phi_{j}(\tau_{t,j})\Bigr)+\Bigl[h(\tau_{1}^{*},\dots,\tau_{J}^{*})-h(\tau_{t,1},\dots,\tau_{t,J})\Bigr].$
By the bounded interaction assumption, we have:
$\Bigl|h(\tau_{1}^{*},\dots,\tau_{J}^{*})-h(\tau_{t,1},\dots,\tau_{t,J})\Bigr|\leq 2\varepsilon.$
Thus, taking expectations and summing over $t=1$ to $K$, the cumulative expected regret satisfies
$\mathrm{Regret}(K)$ $=\sum_{t=1}^{K}\mathbb{E}\left[G(\rho^{*})-G(\rho_{t})\right]$
$\leq\sum_{t=1}^{K}\left[\frac{1}{J}\sum_{j=1}^{J}\,\mathbb{E}\Bigl(\phi_{j}(\tau_{j}^{*})-\phi_{j}(\tau_{t,j})\Bigr)+2\varepsilon\right]$
$=\frac{1}{J}\sum_{j=1}^{J}\sum_{t=1}^{K}\mathbb{E}\left[\phi_{j}(\tau_{j}^{*})-\phi_{j}(\tau_{t,j})\right]+2K\varepsilon.$
By definition, the slot-specific cumulative regret is given by
$\mathrm{Regret}_{j}(K)=\sum_{t=1}^{K}\mathbb{E}\left[\phi_{j}(\tau_{j}^{*})-\phi_{j}(\tau_{t,j})\right],$
so we obtain the final bound:
$\mathrm{Regret}(K)\leq\frac{1}{J}\sum_{j=1}^{J}\,\mathrm{Regret}_{j}(K)+2K\varepsilon.$